# TRiSM for Agentic AI: A Review of Trust, Risk, and Security Management in LLM-based Agentic Multi-Agent Systems

### v2

@@ 9 Discussion
We discuss the broader implications of our findings below:
9.1 Technical Implications of TRiSM for LLM Agent Design
@@
Prior research highlights risks like “excessive agency” [106] where an LLM given too much autonomy or tool access can produce unintended harmful actions (for instance, via hallucination or misinterpreted goals). TRiSM-driven Agentic AI design mitigates these risks by limiting agent autonomy within well-defined safety bounds. Likewise, emerging threats specific to Agentic AI systems such as prompt injection [23] attacks, memory poisoning [137], or cascading hallucinations [217] highlights the need for more built in risk controls. By incorporating anomaly detection and policy-checking modules, an LLM agent can detect deviations from normal behavior and either alert humans or automatically neutralize the threat (e.g. masking a sensitive datum or stopping an unsafe action) .
@@ 9.2 Ethical and Societal Ramifications of Multi-Agent AI
Beyond technical matters, deploying networks of autonomous LLM agents raises pressing ethical and societal questions. Applying TRiSM in this context emphasizes principles of accountability, human oversight, and fairness [169], all of which are vital for public trust in AI systems. Governance in TRiSM recommends that organizations retain clear responsibility for their AI actions, and do not obscure blame behind algorithmic “black boxes”. This implies implementing audit trails and explicable decision logs so that any harmful or biased outcome can be traced and attributed [161]. Recent guidance on trustworthy AI [172] often highlights accountability and explainability as key pillars of trust.
@@
In MAS setups, humans can monitor agent swarms in real time, pause or shut down agents exhibiting anomalies, and adjust policies on the fly [126]. The risk of “user complacency”, trusting an autonomous agent too much, has been noted as a hazard [220]. TRiSM governance counteracts this by formalizing oversight roles and ensuring no AI operates without appropriate human or regulatory supervision.
@@ 9.3 Alignment with Emerging AI Regulations and Standards
The EU AI Act also stresses accuracy, robustness, and cybersecurity for high-risk AI, qualities that TRiSM security management component is designed to ensure (through adversarial resilience, access control, etc.). Similarly, international AI governance standards are emerging that mirror TRiSM’s principles. ISO/IEC 42001:2023, the first global standard for AI management systems, highlights requirements such as transparency, accountability, bias mitigation, safety, and privacy in AI development. By implementing TRiSM, organizations essentially put in place the processes that ISO 42001 [89] and similar standards call for (e.g. leadership oversight, documented risk controls, ongoing monitoring and improvement cycles). Another example is the U.S. NIST AI Risk Management Framework [87], which emphasizes many of the same concepts: identifying risks, embedding governance, and cultivating trustworthiness in AI.
@@ 9.4 Limitations and Current Research Gaps
While the TRiSM-based approach appears promising, we also discuss some limitations and open challenges in current research. First, the classical AI domains (vision, NLP) that have standard test suites, there is no consensus on how to measure an “AI agent” ability to operate safely under TRiSM principles. This makes it difficult to compare different governance strategies or to track progress objectively. We encourage future work to develop evaluation frameworks, possibly extending from adversarial attack simulations [210] or “red-teaming” exercises [221] that can stress-test Agentic AI systems and score their resilience.
Secondly, there is a lack of real-world validation for many TRiSM-inspired controls. Much of the existing literature and tooling for LLM agent safety has been demonstrated in laboratory settings or on narrowly scoped tasks [31]. It remains uncertain how these governance mechanisms perform in complex, open-ended real-world environments.
Another critical gap is adversarial robustness. As we improve defenses, attackers will inevitably adapt [222]. Recent findings show that LLM-based systems remain vulnerable to attacks (for example, hidden prompt injections or subtle data poisoning) that can bypass superficial guardrails [25]. TRiSM solutions today are not foolproof against these. We identify evolving threats and adversarial attacks as ongoing obstacles to trustworthy AI [40].
Finally, organizational and human factors present limitations: implementing TRiSM requires interdisciplinary expertise (AI specialists, security experts, ethicists, legal advisors) and clear governance structures. Many organizations lack the necessary skill sets or frameworks, making TRiSM adoption superficial or inconsistent. Without a strong organizational commitment, even the best technical framework can falter.
@@ 9.5 Future Roadmap for Agentic AI TRiSM
Table 9: Future Roadmap for Agentic AI: Eight Strategic Domains
Scalable Agent Architectures Secure Multi-Agent Collaboration
Explainable Multi-Agent Decisions Human-Centered Trust Design
Autonomous Lifecycle Management Ethics & Governance
Benchmarking & Evaluation Cognitive Capability Expansion
@@
Drawing from our findings and best practices from multiple disciplines, we propose several actionable directions for future research and implementation, as shown in Figure LABEL:fig:roadmap_mindmap. These recommendations span both technical system design improvements and governance-level policy initiatives:
@@
Develop Standardized Evaluation Benchmarks. The community should create open benchmarks and challenge environments to test multi-agent AI governance. For instance, a suite of scenario-based tasks (with built-in threats and ethical dilemmas) could be used to evaluate how well a TRiSM-governed agent system performs relative to one without such controls. This will enable direct comparisons and drive progress on measurable metrics of trust (e.g. frequency of prevented failures or fairness outcomes).
@@
Advance Adversarial Robustness Techniques. Future system design must anticipate a continually evolving threat landscape. Techniques from cybersecurity (e.g. adversarial training, AI model “penetration testing” [223], and formal verification) should be integrated into the LLM agent development pipeline. Cross-disciplinary collaboration with security experts can yield LLM-specific hardening methods, such as dynamic prompt anomaly detectors or robust tool APIs that constrain agent actions. Additionally, creating red-team/blue-team exercises for AMAS , akin to cyber wargames [224] , can help discover vulnerabilities in a controlled way before real adversaries do.
@@
Human-Centered Oversight Tools. We encourage designing better interfaces and protocols for human oversight of Agentic AI systems. Borrowing from human-computer interaction [225] and cognitive engineering [226], researchers could devise dashboards that visualize an agent society’s state, flag important decisions, and allow intuitive human intervention (pausing agents, rolling back actions, etc.).
@@
Regulatory Sandboxes and Compliance-by-Design: Policymakers and industry should collaborate to create regulatory sandboxes for multi-agent AI trials. These would be controlled environments where innovators can deploy Agentic AI systems under supervision, demonstrating TRiSM controls to regulators. Insights from such pilots can inform refinements in both technical standards and regulations. Moreover, adopting a compliance-by-design mindset is key: future AI system designs should bake in the requirements of frameworks like the EU AI Act and ISO 42001 from the start rather than retrofit them.
@@
Cross-Domain Best Practices and Ethical Governance: There is much to learn from other high-stakes domains. For example, the safety engineering field (e.g. aerospace, automotive) has mature practices for redundant controls and failure mode analysis; these could inspire analogous practices in AI agent design. Likewise, ethics boards in biomedical research provide a template for AI ethics committees that review agent behaviors and approve high-risk deployments. We advocate establishing multidisciplinary governance boards that include ethicists, legal experts, domain specialists, and community representatives to oversee significant deployments of autonomous AI.
@@ 10 Conclusion
The rapid advancement of Agentic AI, driven by LLM-based multi-agent systems, marks a transformative shift in autonomous systems, enabling sophisticated collaboration, planning, and decision-making across high-stakes domains such as healthcare, finance, and public services. However, this evolution introduces complex challenges in trust, risk, and security management (TRiSM) that demand a tailored framework to ensure safe, transparent, and accountable operations. This review has systematically explored these challenges, proposing a TRiSM framework structured around four pillars: Governance, Explainability, ModelOps, and Privacy/Security, specifically adapted for LLM-based Agentic Multi-Agent Systems (AMAS).
Key contributions include a novel risk taxonomy capturing unique threats like prompt injection, memory poisoning, and collusive failures, alongside two new metrics: the Component Synergy Score (CSS) and Tool Utilization Efficacy (TUE), which provide practical tools for assessing inter-agent collaboration and tool efficiency. The review also highlights state-of-the-art techniques, such as LIME, SHAP, and decision provenance, to enhance explainability, while advocating for robust security measures like prompt hygiene, sandboxing, and homomorphic encryption to mitigate vulnerabilities.
Looking forward, the proposed research roadmap emphasizes scalable, verifiable, and regulation-compliant Agentic AI systems. Future work should prioritize advancing adversarial robustness, refining governance protocols, and developing standardized benchmarks for trustworthiness and coordination. By aligning Agentic AI development with TRiSM principles, researchers and practitioners can foster systems that are not only innovative but also responsible, ensuring their safe and ethical integration into society.

### v3

@@ 11 Discussion
In this section, we present our findings and outline directions for future work.
11.1 Technical Implications of TRiSM for Agentic AI Design
@@
Prior research highlights risks like “excessive agency” [120] where an LLM given too much autonomy or tool access can produce unintended harmful actions (for instance, via hallucination or misinterpreted goals). TRiSM-driven Agentic AI design mitigates these risks by limiting agent autonomy within well-defined safety bounds. Likewise, emerging threats specific to Agentic AI systems such as prompt injection attacks, memory poisoning , or cascading hallucinations, highlights the need for more built in risk controls. By incorporating anomaly detection and policy-checking modules, an LLM agent can detect deviations from normal behavior and either alert humans or automatically neutralize the threat (e.g. masking a sensitive datum or stopping an unsafe action).
@@ 11.2 Ethical and Societal Ramifications of Multi-Agent AI
Beyond technical matters, deploying networks of autonomous LLM agents raises pressing ethical and societal questions. Applying TRiSM in this context emphasizes principles of accountability, human oversight, and fairness [158], all of which are vital for public trust in AI systems. Governance in TRiSM recommends that organizations retain clear responsibility for their AI actions, and do not obscure blame behind algorithmic “black boxes”. This implies implementing audit trails and explicable decision logs so that any harmful or biased outcome can be traced and attributed . Recent guidance on trustworthy AI [161] often highlights accountability and explainability as key pillars of trust.
@@
In AMAS setups, humans can monitor agent swarms in real time, pause or shut down agents exhibiting anomalies, and adjust policies on the fly . Therisk of “user complacency”, trusting an autonomous agent too much, has been noted as a hazard [193]. TRiSM governance counteracts this by formalizing oversight roles and ensuring no AI operates without appropriate human or regulatory supervision.
@@ 11.3 Alignment with Emerging AI Regulations and Standards
Deploying Agentic AI systems responsibly requires a multifaceted governance approach. By adhering to established regulatory standards (and keeping abreast of new ones), by ensuring auditability for transparency and accountability, and by enforcing policies through technical and organizational means, we can significantly mitigate the risks posed by autonomous AI agents. Moreover, embracing evolving best practices – from governance boards to adaptive, code-driven policies – will help maintain control over these systems as they grow in complexity. The end goal is to harness the benefits of Agentic AI (increased efficiency, creativity, and problem-solving at scale) while staying within legal, ethical, and safe bounds, thereby preserving public trust and meeting compliance obligations.
@@ 11.4 Limitations and Current Research Gaps
While the TRiSM-based approach appears promising, we also discuss some limitations and open challenges in current research. First, the classical AI domains that works on standard test suites, there is no consensus on how to measure an AI Agent vs the Agentic AI ability to operate safely under TRiSM principles. This makes it difficult to compare different governance strategies or to track progress objectively. We encourage future work to develop evaluation frameworks, possibly extending from adversarial attack simulations [194] or “red-teaming” exercises [195] that can stress-test Agentic AI systems and score their resilience.
There is also a lack of real-world validation for many TRiSM-inspired controls. Much of the existing literature and tooling for LLM agent safety has been demonstrated in laboratory settings or on narrowly scoped tasks [25]. It remains uncertain how these governance mechanisms perform in complex, open-ended real-world environments. Another critical gap is adversarial learning. As we improve defenses, attackers will inevitably adapt [196]. Recent findings show that LLM-based systems remain vulnerable to attacks (for example, hidden prompt injections or subtle data poisoning) that can bypass superficial guardrails [19].
Beyond technical aspects, the organizational and human factors present limitations: implementing TRiSM requires interdisciplinary expertise (AI specialists, security experts, ethicists, legal advisors) and clear governance structures. Many organizations lack the necessary skill sets or frameworks, making TRiSM adoption superficial or inconsistent. Without a strong organizational commitment, even the best technical framework can falter.
@@ 11.5 Future Roadmap for Agentic AI TRiSM
Agentic AI Roadmap Scalable Agent Architectures Secure Multi-Agent Collaboration Explainable Multi-Agent Decisions Human-Centered Trust Design Autonomous Lifecycle Management Ethics & Governance Benchmarking & Evaluation Cognitive Capability Expansion
Figure 8: Agentic AI Roadmap: Mindmap Representation
@@
Drawing from our findings and best practices from multiple disciplines, we propose several actionable directions for future research and implementation, as shown in Figure 8. These recommendations span both technical system design improvements and governance-level policy initiatives:
@@
The community should create open benchmarks and challenge environments to test AMAS governance. For instance, a suite of scenario-based tasks (with built-in threats and ethical dilemmas) could be used to evaluate how well a TRiSM-governed Agentic AI systemthat performs relative to one without such controls. This will enable direct comparisons and drive progress on measurable metrics of trust (e.g. frequency of prevented failures or fairness outcomes).
@@
Future system design must anticipate a continually evolving threat landscape. Techniques from cybersecurity (e.g. adversarial training, AI model “penetration testing” [197], and formal verification) should be integrated into the LLM agent development pipeline. Cross-disciplinary collaboration with security experts can yield LLM-specific hardening methods, such as dynamic prompt anomaly detectors or robust tool APIs that constrain agent actions. Additionally, creating red-team/blue-team exercises for AMAS , akin to cyber wargames [198] , can help discover vulnerabilities in a controlled way before real adversaries do.
@@
We encourage designing better interfaces and protocols for human oversight of Agentic AI systems. Borrowing from human-computer interaction [199] and cognitive engineering [200], researchers could devise dashboards that visualize an agent society’s state, flag important decisions, and allow intuitive human intervention (pausing agents, rolling back actions, etc.).
@@
Policymakers and industry should collaborate to create regulatory sandboxes for multi-agent AI trials. These would be controlled environments where innovators can deploy Agentic AI systems under supervision, demonstrating TRiSM controls to regulators. Insights from such pilots can inform refinements in both technical standards and regulations.
@@
There is much to learn from other high-stakes domains. For example, the safety engineering field (e.g. aerospace, automotive) has mature practices for redundant controls and failure mode analysis; these could inspire analogous practices in AI agent design. Likewise, ethics boards in biomedical research provide a template for AI ethics committees that review agent behaviors and approve high-risk deployments. We advocate establishing multidisciplinary governance boards that include ethicists, legal experts, domain specialists, and community representatives to oversee significant deployments of autonomous AI.
@@ 12 Conclusion
The rapid advancement of Agentic AI, driven by LLM-based multi-agent systems, marks a transformative shift in autonomous systems, enabling sophisticated collaboration, planning, and decision-making across high-stakes domains such as healthcare, finance, and public services. However, this evolution introduces complex challenges in trust, risk, and security management (TRiSM) that demand a tailored framework to ensure safe, transparent, and accountable operations. This review has explored these challenges, proposing a TRiSM framework structured around four pillars: Explainability, ModelOps, Security, Privacy and Governance, adapted for LLM-based AMAS. Key contributions include a novel risk taxonomy capturing unique threats like prompt injection, memory poisoning, and collusive failures, alongside two new metrics: the Component Synergy Score (CSS) and Tool Utilization Efficacy (TUE), which provide practical tools for assessing inter-agent collaboration and tool efficiency. The review also highlights state-of-the-art techniques, such as LIME, SHAP, and decision provenance, to enhance explainability, while advocating for robust security measures like prompt hygiene, sandboxing, and homomorphic encryption to mitigate vulnerabilities. Future work should prioritize advancing adversarial robustness, refining governance protocols, and developing standardized benchmarks for trustworthiness and coordination. By aligning Agentic AI development with TRiSM principles, researchers and practitioners can foster systems that are not only innovative but also responsible, ensuring their safe and ethical integration into society.